# Connecting the Dots: A Chain-of-Collaboration Prompting Framework for LLM Agents

### v1

@@ 3.1 Collaborative Knowledge Graph
The construction of the collaborative knowledge graph relies on a large amount of stage-specific knowledge. We adopt the Self-Instruct [55] to generate vertical domain instruction datasets $D_{i}=\{(X_{ij},Y_{ij})\}_{j=1}^{n_{i}}$ for fine-tuning LLMs, where $X_{ij}\in X_{i}$ is the input and $Y_{ij}\in Y_{i}$ is the corresponding output. The datasets are subject to data cleaning and triplet extraction to acquire knowledge for constructing the explicit knowledge graph $\mathcal{KG}_{\text{explicit}}$:
@@
In order to leverage the capabilities of other business workflow agents, we need to identify the tacit variables that play a critical role in the agent’s responses. Since extracting the internal tacit variables of agents is challenging, we focus on the extraction of tacit knowledge from the hidden layers. To achieve this, we introduce counterfactual reasoning. Specifically, we perform causal reasoning, adversarial reasoning, substitution reasoning, extreme counterfactual reasoning, and backward causal reasoning on the questions within the datasets. For each sample $(X_{ij},Y_{ij})$, we generate the corresponding counterfactual input $\tilde{X}_{ij}$:
@@
We then input these questions into the relevant vertical domain agents to generate answers. To represent the output of tacit knowledge influenced by tacit variables, we introduce the tacit variable $\theta_{ij}$, which represents the tacit knowledge or state inside the model and plays a key role in generating counterfactual outputs. The tacit knowledge $h_{ij}$ is the model’s hidden state under the influence of the tacit variable $\theta_{ij}$, which acts as an intermediary in the process of generating counterfactual outputs. The tacit knowledge $h_{ij}$ is influenced by the tacit variable $\theta_{ij}$ where $h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$. The tacit knowledge $h_{ij}$ further influences the generation of the counterfactual output $\tilde{Y}_{ij}$, where $\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The final generation process is expressed as (Appendix A):
[…]
In this process, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ represents the probability of generating the counterfactual output given the tacit knowledge $h_{ij}$ and the counterfactual input $\tilde{X}_{ij}$. $P(h_{ij}\mid\theta_{ij})$ represents the generative distribution of the tacit knowledge $h_{ij}$ given the tacit variable $\theta_{ij}$, and $P(\theta_{ij}\mid\tilde{X}_{ij})$ represents the prior distribution of the tacit variable $\theta_{ij}$ given the counterfactual input $\tilde{X}_{ij}$. Subsequently, these answers are evaluated by a general-purpose LLM, and the evaluation results are classified into three levels: reasonable, ambiguous, and unreasonable. The general-purpose model provides feedback on evaluation opinions and suggestions, which are then used by the vertical domain agent to generate another response. This process iterates until a reasonable answer is obtained.
During knowledge distillation, we apply causal enhancement, with special attention to causal keywords such as “depends on”, “relies on”, and “applies to”. Using the estimated tacit variable $\theta_{ij}$, tacit knowledge is extracted from the counterfactual input-output pairs $(\tilde{X}_{ij},\tilde{Y}_{ij})$ completed through iteration, and a tacit knowledge graph $\mathcal{KG}_{\text{tacit}}$ is constructed:
@@
Finally, by integrating the explicit knowledge graph and the tacit knowledge graph, we construct a collaborative knowledge graph $\mathcal{KG}$ that contains both stage-specific knowledge and inter-stage connectivity:
$\mathcal{KG}=\mathcal{KG}_{\text{explicit}}\cup\mathcal{KG}_{\text{tacit}}$ (5)
@@ 3.2 Causal Chain
To enable the agent to organize the causal relationships between various knowledge stages and generate more accurate and coherent responses, rather than simply listing knowledge, we introduce a causal chain mechanism. When a user submits a request to the agent, the request is first decomposed, and keywords are extracted. These keywords are then input into the knowledge graph to match relevant nodes.
Figure 3: Retrieving relevant knowledge from the collaborative knowledge graph and the process of constructing causal chains.
Specifically, as shown in Figure 3, we use a pre-trained text encoder [45] to convert the extracted keywords and their related nodes into vector representations. For the vector representation $v_{k}$ of keyword $k$ and the vector representation $v_{n}$ of node $n$ in the knowledge graph, we compute the cosine similarity between these vectors to identify the most relevant stage knowledge:
@@
By setting a similarity threshold $\delta$, we select knowledge nodes that are highly related to the semantics of the keywords. For these nodes, we extend their one-hop neighbors to introduce additional associated knowledge and construct a triple-based knowledge structure to enhance the agent’s understanding capability. To address the challenge of cross-stage integration, we adopt a multi-hop neighbor exploration method to build causal chains. By leveraging stage-specific knowledge as a bridge, this approach effectively explains inter-stage knowledge relationships. Specifically, within the stage set $X=\{x_{1},x_{2},x_{3},\dots\}$, for a node $A$ with high semantic similarity, where $A\in x_{1}$, we expand its one-hop neighbor node $B$, where $B\in x_{1}\cap x_{i}$ ($i\neq 1$), and use this as a bridge to explore nodes $C$ from other stages, where $C\in x_{i}$ ($i\neq 1$), to construct the causal chain:
[…]
Using the causal chain mechanism, the agent is able to comprehend and identify the causal relationships between the knowledge we provide. This enables the agent to consider the interdependencies between different components when generating responses to user queries.
Table 1: Performance comparison of different model backbones and multi-agent baselines across three business domains. The performance of our proposed method and all baselines. BERTScore F1 is abbreviated as BS-F. We highlight the best and second-best results.
Backbone Baseline Automotive Pharmaceutical E-commerce
BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L
Qwen2-7B PMC [68] 65.81 13.25 18.37 68.55 17.34 25.93 70.88 24.01 29.91
MedAgents [50] 64.94 12.47 18.34 65.96 10.63 21.13 69.57 17.48 33.62
Debate(short) [20] 65.32 12.66 17.81 70.57 15.78 23.03 72.93 20.85 28.23
Debate(long) [20] 65.56 12.92 18.03 61.14 9.58 12.23 72.89 20.81 28.04
CoA [72] 71.72 22.31 24.01 77.26 29.92 33.70 79.20 38.08 39.02
Cochain 75.39 28.17 28.87 78.48 33.98 35.03 79.91 37.08 40.06
DeepSeek-R1-7B PMC [68] 65.99 11.09 16.00 67.95 23.02 28.18 67.52 16.52 23.53
MedAgents [50] 64.94 12.47 18.34 70.44 24.77 30.73 67.19 14.60 21.18
Debate(short) [20] 66.20 11.45 17.39 71.52 27.03 30.56 70.72 22.92 26.99
Debate(long) [20] 65.94 11.45 17.21 71.77 27.19 30.39 70.68 22.93 27.16
CoA [72] 71.41 12.12 17.55 77.29 38.14 35.43 80.33 43.81 41.54
Cochain 73.19 21.73 24.01 78.97 37.94 36.75 82.68 50.30 47.24
Claude-3.5-haiku PMC [68] 64.11 10.86 10.71 66.51 12.33 13.56 66.96 17.17 16.28
MedAgents [50] 64.45 11.12 11.09 66.23 14.02 14.32 68.11 18.18 16.76
Debate(short) [20] 64.67 11.27 13.44 69.16 14.81 16.06 68.62 19.77 18.25
Debate(long) [20] 65.24 10.88 12.76 67.64 15.32 15.44 68.79 18.06 20.08
CoA [72] 63.39 8.99 10.21 65.22 13.66 14.27 66.92 15.20 15.01
Cochain 69.61 18.73 17.30 72.87 23.59 19.67 71.36 20.15 20.78
@@ 3.3 Prompts Tree
In business workflow tasks, the interconnections between various stages are crucial for effective collaboration. To achieve this, we propose an innovative prompt distillation method aimed at constructing a prompts tree. This method derives stage-specific Q&A from the seed Q&A using a fragmentation approach, where each node at every level of the tree corresponds to prompts from different stages. By querying the prompts tree, it is possible to obtain prompt information that spans the entire business workflow. Specifically, we start with seed Q&A in Stage 1 and utilize the agent from Stage 1 to distill solution-oriented prompts from the answers. The most effective $m$ prompts are selected through self-evaluation. These $m$ prompts are then used to generate new question templates, focusing primarily on how to implement or solve the prompts distilled from the previous stage in the next stage. Each question is answered, forming a new seed Q&A. These Q&A are further involved in the prompt distillation of the current stage, driving the content to extend to the next stage.
In this process, the user need is treated as the root node, with prompts in Stage 1 derived as child nodes based on the user need. These Stage 1 prompts further generate Q&A in Stage 2, from which Stage 2-specific prompts are distilled as child nodes of the Stage 1 prompts. In this way, we continuously build and maintain a prompts tree. By retrieving the prompts tree, we automatically generate a prompt chain that covers the entire business workflow, thus avoiding the difficulties associated with manually constructing prompts. These prompts will guide large language models to generate responses that comprehensively consider the content of each stage, ensuring high collaboration across the stages.
@@ 4 Experiments
In our experiments, we evaluate our collaboration framework and answer the following questions: (1) To what extent does Cochain enhance inter-agent collaboration capabilities? (2) How does Cochain affect agent performance across specialized and cross-domain collaboration tasks? (3) Which key design elements of Cochain are most critical for enhancing collaboration capabilities?
4.1 Experiments Setup
@@
To comprehensively evaluate our method’s impact on business workflow collaboration, we construct novel datasets. Given the scarcity of existing collaborative datasets and their limited ability to reflect complex cooperation patterns in authentic business environments, we develop three business workflow datasets of increasing complexity: e-commerce, pharmaceutical, and automotive workflows. Detailed datasets construction methodology is provided in Appendix B. Notably, the automotive workflow test set underwent rigorous review and refinement by domain experts from a major automotive manufacturer, ensuring its accuracy and practical relevance.
@@
BERTScore [70] utilizes contextual word embeddings to evaluate semantic similarity, while GLEU [62] emphasizes grammatical and semantic quality, and ROUGE-L [35] assesses deep semantic alignment, particularly for extended texts. We employ these metrics as primary indicators of semantic similarity between generated and reference answers. For comprehensive experimental results across additional metrics, refer to Appendix C.
@@
We utilize fifteen LLMs as the backbone of Cochain. With both proprietary and open-source models, we examine the differences between reasoning-capable models (DeepSeek-R1-7B [18]) and standard models (Qwen2-7B [2]). We also investigate using identical versus mixed models (Llama2-7B [52], GLM4-9B, Qwen2-7B, DeepSeek-7B) across workflow stages. All open-source models undergo domain-specific fine-tuning to serve as specialized agents in target domains. For proprietary models, we assess GPT-3.5-turbo, GPT-4o [3], and Claude-3.5-haiku [1] performance.
@@
Our baseline selection encompasses the predominant multi-agent collaboration methodologies (Appendix B). Specifically, PMC [68] implements multi-agent collaboration through Hierarchical Planning. MedAgents [50] facilitates collaborative decision-making via discussion and voting mechanisms. Debate [20] enables collaboration through argumentative discourse, modulating agent confidence through varying debate durations. CoA [72] establishes chain-based collaboration where agents interact through dedicated CU. Besides, we compare the performance of single-agent prompt engineering techniques, namely CoT [57] and ToT [66].
Figure 4: (a) Expert evaluation & (b) GPT-4 scores.
Figure 5: Reasoning time differential.
Backbone Method BS-F GLEU ROUGE-L
GPT-3.5-turbo + IO 72.78 15.28 22.06
+ COT [57] 71.28 10.91 17.73
+ TOT [66] 66.79 10.15 17.92
+ Cochain 76.92 28.68 30.73
GPT-4o + IO 71.69 22.71 23.48
+ COT [57] 72.70 24.03 24.44
+ TOT [66] 66.29 12.54 17.82
+ Cochain 74.51 25.17 25.22
Table 3: Result of different skip stages on evaluation metrics.
@@
4.2 Enhancing Collaborative Reasoning Capabilities in LLM Agents with Cochain
Overall Results of Cochain.
Table 1 demonstrates Cochain’s performance on business workflow tasks across three diverse language models. Results show Cochain consistently outperforms all baselines by significant margins across all datasets. Cochain also exhibits substantial advantages in computational efficiency, with shorter reasoning times, fewer output tokens, and lower operational costs compared to other baselines (Appendix C.1). Figure 7 reports consistent performance gains achieved by Cochain as model parameters increase, demonstrating its strong scalability potential with more powerful foundation models. As shown in Figure 5, the reasoning time differential between thinking and non-thinking models across increasing reasoning rounds reveals that with PMC, this differential grows dramatically as rounds increase, while with Cochain, it remains relatively moderate. This demonstrates Cochain’s superior efficiency in handling complex reasoning tasks, with its performance advantage becoming increasingly pronounced as reasoning complexity increases.
Comparative Analysis of Cochain Against Prompting Engineering Approaches.
Table 3 presents a comparative evaluation of Cochain against established prompting methodologies. The results show that while CoT and ToT approaches enhance reasoning capabilities, they exhibit limited collaborative advantages compared to our framework. Specifically, these traditional techniques focus on isolated reasoning paths, whereas Cochain enables agents to incorporate both upstream and downstream factors into decision-making, yielding more comprehensive solutions for complex, interdependent problems across multiple domains of expertise.
Qualitative Evaluation.
Given the absence of established metrics for evaluating LLM agent collaboration, we conduct a two-fold validation approach. First, we recruit five experts from a major automotive manufacturer for evaluation. These experts judge Qwen2-7B with Cochain to outperform GPT-4 in approximately 60% of cases (Figure 5(a)). We also employ GPT-4 for quantitative assessment across five scoring rounds, comparing performance with and without fine-tuning and our framework. Figure 5(b) shows models enhanced with Cochain (Raw* and Finetune*) demonstrate significantly higher median scores, with the fine-tuned version achieving nearly optimal performance (median >9.7). These evaluations provide compelling evidence for Cochain’s real-world effectiveness.
Figure 6: Parameter analysis results. In the first five plots, the top-n value is fixed, while in the last plot, the number of prompts is fixed.

### v2

@@ 3.1 Collaborative Knowledge Graph
Building a collaborative knowledge graph requires abundant stage-specific knowledge. We obtain it from the training sets of the workflow datasets, denoted as $D_{i}=\{(X_{ij},Y_{ij})\}_{j=1}^{n_{i}}$, where $i\in\{1,\dots,N\}$ indexes workflow stages, $N$ is the number of stages, and $j\in\{1,\dots,n_{i}\}$ indexes the $n_{i}$ training instances in stage $i$. Here $X_{ij}\in X_{i}$ is the input and $Y_{ij}\in Y_{i}$ is the corresponding output. After data cleaning and triplet extraction, we construct an explicit knowledge graph $\mathcal{KG}_{\text{explicit}}$:
@@
Business workflows also contain tacit knowledge that is rarely stated in the original datasets, such as feasibility constraints, implementation caveats, and stage-specific heuristics. Since such considerations are not explicitly available, we elicit tacit workflow knowledge from the agent’s responses by probing the agent under counterfactual perturbations and attributing response variations to unobserved internal considerations.
For each sample $(X_{ij},Y_{ij})$, we generate a counterfactual input $\tilde{X}_{ij}$ by editing $X_{ij}$ to perturb one salient condition while preserving intent, using templates for causal, adversarial, substitution, extreme, and backward-causal variants:
@@
In our setting, a vertical domain agent is the stage-specialized LLM assigned to stage $i$ (fine-tuned on stage-$i$ data for open-source backbones). We query this agent with $\tilde{X}_{ij}$ to obtain the counterfactual output $\tilde{Y}_{ij}$. To capture the unobserved considerations that shape how the agent resolves $\tilde{X}_{ij}$, we introduce a latent variable $\theta_{ij}$ and an intermediate latent state $h_{ij}$ influenced by $\theta_{ij}$, where $h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$, and $\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The resulting generation process can be written as:
$\begin{split}P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})&=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})\\ &\quad\cdot P(h_{ij}\mid\theta_{ij})\,P(\theta_{ij}\mid\tilde{X}_{ij})\,d\theta_{ij}\end{split}$ (3)
Here, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ is the likelihood of generating the counterfactual output given $h_{ij}$ and $\tilde{X}_{ij}$, $P(h_{ij}\mid\theta_{ij})$ describes how tacit considerations shape the latent state, and $P(\theta_{ij}\mid\tilde{X}_{ij})$ is a prior over tacit considerations conditioned on the counterfactual input.
To improve reliability, each counterfactual answer is evaluated by a general-purpose LLM and assigned one of three labels: reasonable, ambiguous, or unreasonable. The evaluator also provides concise feedback and revision suggestions, which the vertical agent uses to regenerate a response. This loop repeats until a reasonable answer is obtained. During knowledge distillation, we apply causal enhancement by emphasizing causal cues such as “depends on”, “relies on”, and “applies to”, which helps preserve cross-stage dependency signals. Using the counterfactual input-output pairs $(\tilde{X}_{ij},\tilde{Y}_{ij})$ produced by the refinement loop, we extract tacit triples and build a tacit knowledge graph $\mathcal{KG}_{\text{tacit}}$:
@@
Finally, we integrate explicit and tacit knowledge to form the collaborative knowledge graph $\mathcal{KG}$, which captures both stage-specific knowledge and inter-stage connectivity:
$\mathcal{CKG}=\mathcal{KG}_{\text{explicit}}\cup\mathcal{KG}_{\text{tacit}}$ (5)
@@ 3.2 Causal Chain
To help the agent organize cross-stage dependencies instead of merely listing retrieved facts, we introduce a causal chain mechanism. Given a user request, we decompose it and extract keywords, which are then used to retrieve relevant nodes from the collaborative knowledge graph.
Following Figure 2, we use a pre-trained text encoder Reimers and Gurevych (2019) to embed each keyword and candidate knowledge node. For the embedding $v_{k}$ of keyword $k$ and the embedding $v_{n}$ of node $n$, we compute cosine similarity:
@@
With a similarity threshold $\delta$, we select nodes that are semantically related to the query. For each selected node, we expand one-hop neighbors to introduce closely associated evidence and form a local triple-based structure. To connect knowledge across workflow stages, we further perform multi-hop exploration through bridge entities. Let the workflow stages be $\mathcal{S}=\{s_{1},s_{2},s_{3},\dots\}$. For a high-similarity node $A$ belonging to stage $s_{1}$, we expand a neighbor $B$ that appears in both $s_{1}$ and some other stage $s_{i}$ with $i\neq 1$. We then use $B$ as a bridge to reach a node $C$ in stage $s_{i}$, forming a cross-stage causal chain:
$\text{CausalChain}(A,C)=A\xrightarrow{B}C$ (7)
The causal chain provides a compact, stage-connected context that highlights how constraints and decisions propagate across stages, improving coherence and feasibility in the final response.
Backbone Method BS-F GLEU ROUGE-L
GPT-3.5-turbo + IO ${\text{72.84}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{15.09}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{21.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$
+ CoT ${\text{71.61}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{11.15}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{17.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.16}}$
+ ToT ${\text{66.69}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{10.08}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.04}}$ ${\text{17.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$
+ Cochain ${\text{76.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.03}}$ ${\text{28.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.12}}$ ${\text{30.65}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$
GPT-4o + IO ${\text{71.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{23.05}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{23.79}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$
+ CoT ${\text{72.57}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.15}}$ ${\text{23.76}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{24.15}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$
+ ToT ${\text{66.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{12.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{17.52}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$
+ Cochain ${\text{74.46}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{25.16}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{25.07}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$
Baseline Backbone
DeepSeek-V3.1 Claude-3.5-haiku
PMC 0.12 0.29
MedAgents 0.07 0.26
Debate 0.32 0.31
CoA 0.27 0.37
Cochain 0.33 0.38
Table 3: Accuracy comparison on Auto-SLURP benchmark.
@@ 3.3 Prompts Tree
In business workflow tasks, effective collaboration depends on whether stage-specific reasoning can be carried forward in a consistent and reusable manner. We propose a prompt distillation procedure that offline constructs a Prompts Tree, which stores reusable, solution-oriented prompts and explicitly encodes cross-stage dependencies. We model the Prompts Tree as a rooted tree $T=(V,E)$. Each node $v\in V$ is associated with a stage label $s(v)\in\mathcal{S}$, a distilled prompt $p(v)$, and its supporting evidence (the Q&A from which the prompt is distilled). Each directed edge $(u\!\rightarrow\!v)\in E$ represents a dependency where prompt $p(u)$ from an upstream stage induces a downstream question template, whose answer leads to prompt $p(v)$ at a later stage. Querying $T$ returns a prompt chain along a root-to-leaf path, providing structured, cross-stage guidance during inference.
Tree construction. We start from a seed Q&A at Stage 1. Given the Stage 1 answer, the Stage 1 agent distills a set of candidate prompts that are solution-oriented and constraint-aware, focusing on actionable decisions, feasibility conditions, and implementation requirements. Each candidate prompt is then self-evaluated for utility and executability, and we retain the top $m$ prompts. For each retained prompt, we generate a downstream question template that asks how to implement, validate, or satisfy this prompt in the next stage. The next-stage agent answers these questions, producing new Q&A pairs that serve as seeds for prompt distillation at the next stage. This distill–self-evaluate–generate-next-question–expand loop repeats across stages, causing the tree to grow as upstream prompts branch into multiple downstream realizations.
Node and edge semantics. The user need is treated as the root node. Prompts distilled at Stage 1 become children of the root. For a node $u$ at stage $s_{k}$, each outgoing edge corresponds to a generated question template $q_{u\rightarrow}$ targeting stage $s_{k+1}$. The resulting answer yields a child node $v$ at stage $s_{k+1}$ with distilled prompt $p(v)$. This construction explicitly records cross-stage dependencies, since each child prompt is conditioned on satisfying or operationalizing its parent prompt under downstream constraints.
Retrieval and usage. Inspired by how the human brain integrates fragmented knowledge into coherent thought Courellis et al. (2024), the Prompts Tree organizes fragmented stage-specific guidance into an ordered chain. At inference time, we retrieve a prompt chain $\pi=(v_{0}\!\rightarrow v_{1}\!\rightarrow\cdots\!\rightarrow v_{L})$ that best matches the input query, and use the prompts $\{p(v_{\ell})\}_{\ell=0}^{L}$ as structured guidance so the backbone LLM reasons through the workflow stages in a consistent order. This avoids brittle manual prompt construction while improving cross-stage coordination with low token overhead.
@@ 4 Experiments
We study when collaboration helps and when it hurts in multi-stage business workflows, and evaluate whether Cochain provides a practical middle ground via reusable collaboration artifacts. Accordingly, we benchmark overall performance against strong baselines, analyze key design factors and hyperparameters, compare jointly fine-tuned and stage-wise specialized deployment regimes, and report efficiency and cost.
4.1 Experimental Setup
@@
We evaluate Cochain on two benchmarks across four datasets: Auto-SLURP Shen and Shen (2025), which assesses end-to-end multi-stage workflows, and MSCoRe Lei et al. (2025), which focuses on multi-stage reasoning and collaboration. The automotive workflow test set is reviewed and refined by domain experts from a major automobile manufacturer to ensure practical industry relevance.
@@
We report BERTScore Zhang et al. (2019), GLEU Wu et al. (2016), and ROUGE-L Lin (2004) to measure semantic similarity and generation quality between model outputs and reference answers.
@@
We use 17 backbone LLMs, covering both proprietary and open-source models. We compare reasoning-capable models (e.g., DeepSeek-R1-7B DeepSeek-AI (2025)) with standard models (e.g., Pangu-1B Chen et al. (2025), Qwen2-7B 29). We also study using identical backbones across stages versus mixed backbones (Llama2-7B Touvron et al. (2023b), GLM4-9B, Qwen2-7B, DeepSeek-7B) for stage specialization. For open-source backbones, we perform stage-wise fine-tuning on MSCoRe: the agent at stage $k$ is fine-tuned using only the training split of stage-$k$ data, without using any validation/test instances. Proprietary models include GPT-3.5-turbo, GPT-4o Achiam et al. (2023), DeepSeek-V3.1 DeepSeek-AI (2024), and Claude-3.5-haiku.
Figure 3: Expert evaluation and GPT-4 scoring results. Panel (a) compares Qwen2-7B augmented with Cochain against GPT-4 and expert-written references (100 cases). In panel (b), the asterisk indicates augmentation with Cochain.
@@
Our baseline selection encompasses the predominant multi-agent collaboration methodologies. Specifically, PMC Zhang et al. (2025a) implements multi-agent collaboration through hierarchical planning. MedAgents Tang et al. (2024) facilitates collaborative decision-making via discussion and voting mechanisms. Debate Du et al. (2024) enables collaboration through argumentative discourse. CoA Zhang et al. (2024b) establishes chain-based collaboration. Besides, we compare the performance of single-agent prompt engineering methods, namely CoT Wei et al. (2022) and ToT Yao et al. (2024).
@@
Table 4: Result of different skip stages on evaluation metrics.
4.2 Result
Overall Results.
Table 1 summarizes performance across three business domains and four backbones. The results support our central claim that effective workflow collaboration requires balancing Insufficient and Excessive Collaboration. Compared with multi-agent baselines that rely on rich interaction (often leading to excessive collaboration and diluted decisions), Cochain consistently achieves higher quality across domains. Meanwhile, Cochain also improves substantially on a small backbone, indicating that it mitigates insufficient collaboration by injecting stage-aware constraints and dependencies through reusable artifacts rather than relying on stronger backbones alone. For instance, with Pangu-1B on Automotive, Cochain raises BS-F from 72.28 (PMC) to 74.35 and improves ROUGE-L from 22.89 to 26.01, suggesting that the framework helps the agent stay constraint-aware while remaining focused on the core decision.
Benchmark Transfer on Auto-SLURP.
Table 3 confirms the advantage on an end-to-end workflow benchmark. Cochain achieves the highest accuracy under both DeepSeek-V3.1 and Claude-3.5-haiku backbones, reaching 0.33 and 0.38, respectively. This result supports that Cochain generalizes beyond MSCoRe-style text similarity metrics and remains effective when collaboration must be executed across workflow stages.
Comparison to Prompting Baselines.
Table 2 compares Cochain with IO, CoT, and ToT on proprietary backbones. While CoT and ToT enhance reasoning in a single-agent setting, they provide limited support for carrying stage constraints across a workflow. Cochain yields stronger generation quality under both GPT-3.5-turbo and GPT-4o. On GPT-3.5-turbo, Cochain increases BS-F from 72.84 (IO) to 76.94 and improves ROUGE-L from 21.98 to 30.65, demonstrating that the benefits go beyond deeper single-path reasoning.
Expert Evaluation and GPT-4 Scoring.
Figure 3 provides external validation beyond automatic metrics. In panel (a), domain experts judge Cochain-augmented answers to be better in 40 cases and worse in 9 cases, with 51 cases considered similar. Panel (b) shows higher GPT-4 scores for settings augmented with Cochain, and the fine-tuned augmented setting reaches near-ceiling scores, corroborating the expert preference trend.
Figure 4: Hyperparameter influence of prompt budget and retrieval size. (a) Fix top-$n$ retrieved nodes and vary prompt count. (b) Fix prompt count and vary top-$n$.